How much 3.5% milk and 1.0% milk must be mixed to produce 220 gallons of 2.0% milk?
Answer by checkley77(12844)

3.5%=.035 TIMES SOME NUMBER OF GALLONS PLUS 1%=.01 TIMES THE REST OF THE GALLONS (220-X) RESULTS IN THE 2% OR .O2 TIMES THE ENTIRE 220 GALLONS.
----------------------------------------------------------------------------
.035x+.01(220-x)=.02*220
.035x+2.2-.01x=4.4
.025x=4.4-2.2 (.035x-.010x=.025x)
.025x=2.2
x=2.2/.025
x=88 gallons of 3.5% milk is used.
220-88=132 gallons of 1% is used.
Proof:
.035*88+.01*132=.02*220
3.08+1.32=4.4
4.4=4.4